OpenStudy (amonoconnor):

A question pertaining to Sigma Notation! When evaluating a sum, if the part to the right of the sigma is "k^3 - 45", can it be analyzed, using summation rules, as follows? -> "((n(n+1))/2)^2 + n(45)" ?? Any and all help is greatly appreciated! :)

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^n (k^3 - 45)\)

OpenStudy (anonymous):

Assuming it's a finite sum, sure. You have to be careful if you're summing an infinite series, though...

OpenStudy (amonoconnor):

Thank you so much! I just wanted to double check my conceptual understanding there! :)

OpenStudy (amonoconnor):

Yes, there is a finite limit:)

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^n (k^3 - 45) = \sum \limits_{k=1}^n (k^3 ) - \sum \limits_{k=1}^n ( 45) \)

ganeshie8 (ganeshie8):

u wlc :)

OpenStudy (amonoconnor):

So, say n was 5, and k was 1.... The second part is still just going to be found as 45? Not 5(45)?

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^5 (k^3 - 45) = \sum \limits_{k=1}^5 (k^3 ) - \sum \limits_{k=1}^5 ( 45) \) \(\large \qquad \qquad~~~~~ = \sum \limits_{k=1}^5 (k^3 ) - 45\sum \limits_{k=1}^5 ( 1) \)

OpenStudy (amonoconnor):

... And the second part (aside from the coefficient of 45 now), is equal to five?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^5 ( 1) = 1 + 1 + 1 + 1 + 1 = 5\)

OpenStudy (amonoconnor):

...how? .. If there 's no place to plug in values of k? Does this method/ rule hold true for all sigmas with an integer to the right of the sigma, but no variable?

OpenStudy (amonoconnor):

I see what happened, I just don't get how that... Happens, If there is not a place for k values...

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^5 ( 1) = 1 + 1 + 1 + 1 + 1 = 5\) is same as \(\large \sum \limits_{k=1}^5 ( k^0 1) = (1^0)1 + (2^0)1 + (3^0)1 + (4^0)1 + (5^0)1 = 5\)

OpenStudy (amonoconnor):

..... Ahhh, lightbulb!!! :D (the world needs more people like you ganeshie8)

ganeshie8 (ganeshie8):

the definition of SUM notation is :- u simply take the sums of the right side thing, each time changing the value of variable \(k\)

ganeshie8 (ganeshie8):

lol, okay so below just some examples :- \(\large \sum \limits_{k=1}^5 ( k) = 1 + 2 + 3 +4 + 5 \)

ganeshie8 (ganeshie8):

\(\large \sum \limits_{k=1}^5 ( k^2) = 1^2 + 2^2 + 3^2 +4^2 + 5^2 \)

ganeshie8 (ganeshie8):

\( \sum \limits_{k=1}^5 ( k^2(k-k^{300})) = \\ \\ 1^2(1-1^{300}) + 2^2(2-2^{300}) + 3^2(3-3^{300}) +4^2(4-4^{300}) + 5^2(5-5^{300}) \)

ganeshie8 (ganeshie8):

\(\sum \) is just a notational thing, how u evaluate the sums is entirely a different thing.... u could use sum of first n terms formula or sum of squares of first n terms formula... or watever u r familiar with

OpenStudy (amonoconnor):

More complex, but I still follow:) it was mainly the lack of a K to the right of the sign that was confusing me:) but I feel totally confident in how to navigate it now! Thank you again! :)

ganeshie8 (ganeshie8):

np :)